A resonant circuit in a radio receiver is tuned to a certain station when the inductor has a value of 0.214 mH and the capacitor has a value of 21.2 pF. Calculate the frequency of the radio station Submit Answar Tries 0/12 What is the wavelength of the radiowaves emited by the radio station? Submlt Answar Tries 0/12
Answer #1

a) Capactive reactance = inductive reactance

1/2pi*f*C = 2pi*f*L

1/2pi*f * 21.2*10^-12 = 2pi*f*0.214*10^-3

f = 2.36*10^6 Hz

b) wavelength = c/f

= 3*10^8 / 2.36*10^6 = 126.96 m

= 127 m